# F-Actor: Controllable Conversational Behaviour in Full-Duplex Models

### 05 Mar 2026 v2

@@ Abstract
Spoken conversational systems require more than accurate speech generation to have human-like conversations: to feel natural and engaging, they must produce conversational behaviour that adapts dynamically to the context. Current spoken conversational systems, however, rarely allow such customization, limiting their naturalness and usability. In this work, we present the first open, instruction-following full-duplex conversational speech model that can be trained efficiently under typical academic resource constraints. By keeping the audio encoder frozen and finetuning only the language model, our model requires just 2,000 hours of data, without relying on large-scale pretraining or multi-stage optimization. The model can follow explicit instructions to control speaker voice, conversation topic, conversational behaviour (e.g., backchanneling and interruptions), and dialogue initiation. We propose a single-stage training protocol and systematically analyze design choices. Both the model and training code will be released to enable reproducible research on controllable full-duplex speech systems. [^1] https://github.com/MaikeZuefle/f-actor
@@ 1 Introduction
Recent work has begun to address full-duplex modelling using a variety of architectural choices: training special predictors for overlapping speech (Ruede et al., 2017; Chen et al., 2025b), modelling user and system speech as separate streams (défossez2024moshispeechtextfoundationmodel; Hu et al., 2025) or by interleaving user and system chunks in a single sequence (Veluri et al., 2024; Lee et al., 2025). While the latter approaches enable models to handle interruptions and backchannels, they typically do not model them explicitly on the system side. That is, models are trained to robustly handle overlapping speech produced by the user, but are rarely studied in terms of whether the system itself interrupts or backchannels. As a result, existing evaluations focus on reactive rather than proactive conversational behaviour (Peng et al., 2025; Lin et al., 2025).
@@
    We systematically experiment with and analyze different design and architectural choices.
@@ 2 Background
Full-Duplex Modelling. Full-duplex interaction, simultaneous speaking and listening, is often modelled using parallel user and system speech streams (Hu et al., 2025; défossez2024moshispeechtextfoundationmodel). To enable the backbone LLMs to process speech streams, raw audio is typically mapped to discrete acoustic units (DAUs), which can be treated as tokens in the model vocabulary. Each speech stream then emits DAUs at every timestep, including silence DAUs, and embeddings for each stream are combined by summation (défossez2024moshispeechtextfoundationmodel; Hu et al., 2025), or by token-wise fusion (Wang et al., 2025) to allow concurrent processing. To produce speech, the generated DAUs are mapped back into speech waveforms with pre-trained vocoders.
@@
Discrete Acoustic Units. DAUs are typically produced by neural audio codecs such as Mimi (défossez2024moshispeechtextfoundationmodel), EnCodec (Défossez et al., 2022), or SoundStream (Zeghidour et al., 2021). These codecs compress audio into discrete codes optimized for reconstruction and are widely used as both inputs and outputs in recent speech language models (défossez2024moshispeechtextfoundationmodel; Hu et al., 2025).
The codec models represent each audio frame using multiple discrete codes, drawn from multiple codebooks. The choice of quantization determines how these codes are modelled. Dependent codebooks, typically produced by residual vector quantization (Zeghidour et al., 2021, RVQ), require hierarchical prediction: the language model predicts the first code, while additional modules generate the remaining layers (défossez2024moshispeechtextfoundationmodel; Wang et al., 2023). In this work, we choose independent codebooks from finite scalar quantization (Mentzer et al., 2024, FSQ), which simplify modelling by allowing all codes for a frame to be predicted jointly (Hu et al., 2025). Additional details on encoding and reconstruction are provided in App.˜A.
@@ 3 Related Work
End-to-end architectures typically support these behaviours natively and have therefore become increasingly popular. Among them, Moshi (défossez2024moshispeechtextfoundationmodel) and its variants (Ohashi et al., 2025; Shi et al., 2025) represent a common design choice: employing separate system, user, and text streams, and codec-based models for speech input and output (défossez2024moshispeechtextfoundationmodel; Défossez et al., 2022; Zeghidour et al., 2021). Other approaches use continuous features at the input (Hu et al., 2025; Fu et al., 2025), or at both input and output (Yu et al., 2024). Some architectures avoid multiple input–output streams by interleaving modalities (Veluri et al., 2024; Lee et al., 2025; Yu et al., 2024), or by embedding them jointly at the token level (Wang et al., 2025). For more details, we refer the reader to Arora et al. (2025a).
@@ 4 Model
F-Actor, our full-duplex model, is based around an LLM backbone, augmented with a speaker embedding model and an audio encoder. Unlike previous work (défossez2024moshispeechtextfoundationmodel; Hu et al., 2025), F-Actor is designed with instruction following in mind, that is, incorporating both a speaker embedding and an instruction-following prompt. In the following, we describe each component of the system in detail. An overview is provided in Fig.˜1.
@@
NanoCodec has two advantages: (i) it uses FSQ quantization (Mentzer et al., 2024) so the four codebooks are independent, enabling parallel prediction without requiring a depth-transformer architecture (défossez2024moshispeechtextfoundationmodel), and (ii) it has shown strong performance on speech synthesis (Hu et al., 2025).
@@
Text Stream. Following Hu et al. (2025); défossez2024moshispeechtextfoundationmodel, we not only generate audio, but also experiment with generating the corresponding text, using an additional system-side text stream. To do so, we need to align the text tokens and audio DAUs (alignment strategies are described in Section˜6). After embedding the text stream, we add it to the system’s speech stream. The LLM uses this summed embedding as input. We then use the LLM’s original language-modeling head to predict the next text token.
@@ 6.1 Data
Text-Speech Alignment. As described in Section˜4, we experiment with generating not only speech, but also the corresponding text. We explore both word- and utterance-level alignment of speech and text, as prior work has shown mixed results (défossez2024moshispeechtextfoundationmodel; Hu et al., 2025). To this end, we align text and speech tokens using forced alignment with Kaldi (Povey et al., 2011) and a public model888[Uncaptioned image] kaldi-model-m13 trained on Librispeech Panayotov et al. (2015). We exclude conversations with text–speech mismatches due to occasional TTS deviations from the transcript, resulting in 74.4% of the original training set.
@@ 6.2 Training Details
To maximize efficiency, our approach utilizes a single training stage that avoids additional speech or text pretraining as required by prior models such as défossez2024moshispeechtextfoundationmodel. Only the LLM backbone and the linear DAU heads are trained, while the speech encoder remains frozen. Training is performed on four NVIDIA A100-SXM4-40GB GPUs with early stopping based on validation performance, resulting in around 48 hours of training per model. We use a maximum sequence length of 2,048 tokens, a per-GPU batch size of 1, and 8 gradient accumulation steps. Training and inference parameters are provided in Tab.˜6 in App.˜D.
@@ 6.3 Experiments
Text-Stream. We examine the effect of adding a text stream on the system side (see Section˜4), enabling the model to predict not only audio but also the corresponding text (défossez2024moshispeechtextfoundationmodel; Hu et al., 2025). We further experiment with introducing special tokens that precede BCs and interruptions (BC/I tok.) to help the model generate the numbers specified in the prompt.
@@
Audio Delay. We experiment with delaying the audio stream by up to two tokens relative to the text stream, generating text tokens first while padding the audio stream before audio token generation begins (défossez2024moshispeechtextfoundationmodel; Hu et al., 2025).
RVQ vs. FSQ. All primary experiments use an FSQ encoder to benefit from independent codebooks. To test the generality of our multi-head DAU prediction architecture, we also evaluate an RVQ-based setup using Mimi101010[Uncaptioned image] kyutai/mimi (défossez2024moshispeechtextfoundationmodel) with 8 codebooks. This allows us to verify whether predicting DAUs for each codebook in parallel remains effective with RVQ quantization.
@@ 7.3 Comparing Turn-Taking to SOTA Models
In Tab.˜4, we compare turn-taking behaviour to Moshi (défossez2024moshispeechtextfoundationmodel) and dGSLM (Nguyen et al., 2023), with all metrics reported as cumulative durations per category. For all four metrics, Inter-Pausal Units (IPUs), intra-turn Pause, between-turn Gap and Overlap, our model exhibits behavior similar to the training data and the baseline models. Like Moshi and non-cascaded dGLSM, our model’s pauses are generally longer than the gaps, reflecting real-world conversational statistics (Heldner and Edlund, 2010).
@@
7.3.1 The Role of Sampling Parameters
Similar to défossez2024moshispeechtextfoundationmodel, we observe that sampling temperature has a significant impact on model behaviour. Results for different temperatures (0.60.61.01.0) using our best-performing model (word-level alignment; audio delay of 2) on the dev set are reported in Tab.˜7 in App.˜E. As temperature increases, the dialogue becomes more dynamic, with more BCs and interruptions, more balanced speaking time between the speakers, and improved coherence and adherence to the narrative. However, higher temperatures also lead to slightly lower UTMOS scores and higher WER, reflecting a mild impact on speech quality. In our setup, a temperature of $0.9$ provides the best trade-off, and all results reported in this paper use this setting.
@@ A.1 Discrete Representations of Speech
The first category of discrete representations fo speech consists of units often termed “acoustic tokens” or Discrete Acoustic Units (DAUs), derived from reconstruction-based Neural Audio Codecs (NACs) such as Mimi, EnCodec, or SoundStream (défossez2024moshispeechtextfoundationmodel; Défossez et al., 2022; Zeghidour et al., 2021). These units are optimized to compress and perfectly reconstruct the input signal. Consequently, they encode all acoustic information, including speaker identity, prosody, and recording conditions, into a single stream that can be decoded directly back into a waveform. While this enables end-to-end modeling such as in Moshi and VALL-E (défossez2024moshispeechtextfoundationmodel; Wang et al., 2023), the high information density of DAUs can make modeling stability challenging.
@@ A.2 Quantization Methods
The choice of quantization method fundamentally dictates the architecture of the Speech Language Model, specifically regarding how it models the multiple codes representing a single time step. The two primary paradigms involve dependent codebooks, typically resulting from Residual Vector Quantization (Zeghidour et al., 2021, RVQ), and independent codebooks, such as those derived from Finite Scalar Quantization (Mentzer et al., 2024, FSQ). In RVQ-based models like EnCodec or Mimi (Défossez et al., 2022; défossez2024moshispeechtextfoundationmodel), the codebooks operate sequentially where each quantizes the residual error of the previous one. This creates a strong dependency among the codes within a single frame. Therefore, generation often requires a hierarchical approach. The primary language model autoregressively predicts the code for the first RVQ layer, while a secondary mechanism, such as the Depth Transformer in Moshi or the nonautoregressive model in VALL-E, predicts the codes for subsequent RVQ layers for that frame (défossez2024moshispeechtextfoundationmodel; Wang et al., 2023). Conversely, methods like NanoCodec use FSQ to project the latent space into independent subspaces, resulting in codebooks that are statistically independent (Casanova et al., 2025). This independence simplifies the modeling task by allowing the Language Model to predict all codes for a frame simultaneously. Such models, as seen in Hu et al. (2025), typically utilize a multiple head output layer, termed a Hydra head, which predicts the indices for all independent codebooks in parallel without the need for an additional sequential depthwise modeling step.
@@ Appendix D Training Details
For inference, we show that sampling parameters, in particular the temperature, have a great influence on model performance in Section˜7.3.1. We find that for our setup, a temperature of $0.9$ works best and hence report all results (except for those in Tab.˜7) with this temperature.

### 15 Apr 2026 v3

@@ Abstract
Spoken conversational systems require more than accurate speech generation to have human-like conversations: to feel natural and engaging, they must produce conversational behaviour that adapts dynamically to the context. Current spoken conversational systems, however, rarely allow such customization, limiting their naturalness and usability. In this work, we present the first open, instruction-following full-duplex conversational speech model that can be trained efficiently under typical academic resource constraints. By keeping the audio encoder frozen and finetuning only the language model, our model requires just 2,000 hours of data, without relying on large-scale pretraining or multi-stage optimization. The model can follow explicit instructions to control speaker voice, conversation topic, conversational behaviour (e.g., backchanneling and interruptions), and dialogue initiation. We propose a single-stage training protocol and systematically analyze design choices. Both the model and training code is released to enable reproducible research on controllable full-duplex speech systems. [^1] https://github.com/MaikeZuefle/f-actor
@@ 1 Introduction
Recent work has begun to address full-duplex modelling using a variety of architectural choices: training special predictors for overlapping speech (Ruede et al., 2017; Chen et al., 2025b), modelling user and system speech as separate streams (Défossez et al., 2024; Hu et al., 2025) or by interleaving user and system chunks in a single sequence (Veluri et al., 2024; Lee et al., 2025). While the latter approaches enable models to handle interruptions and backchannels, they typically do not model them explicitly on the system side. That is, models are trained to robustly handle overlapping speech produced by the user, but are rarely studied in terms of whether the system itself interrupts or backchannels. As a result, existing evaluations focus on reactive rather than proactive conversational behaviour (Peng et al., 2025; Lin et al., 2025).
@@
    We systematically experiment with and analyze different design and architectural choices, such as RVQ and FSQ encoders, text alignment variants, audio delays or special tokens.
@@ 2 Background
Full-Duplex Modelling. Full-duplex interaction, simultaneous speaking and listening, is often modelled using parallel user and system speech streams (Hu et al., 2025; Défossez et al., 2024). To enable the backbone LLMs to process speech streams, raw audio is typically mapped to discrete acoustic units (DAUs), which can be treated as tokens in the model vocabulary. Each speech stream then emits DAUs at every timestep, including silence DAUs, and embeddings for each stream are combined by summation (Défossez et al., 2024; Hu et al., 2025), or by token-wise fusion (Wang et al., 2025) to allow concurrent processing. To produce speech, the generated DAUs are mapped back into speech waveforms with pre-trained vocoders.
@@
Discrete Acoustic Units. DAUs are typically produced by neural audio codecs such as Mimi (Défossez et al., 2024), EnCodec (Défossez et al., 2022), or SoundStream (Zeghidour et al., 2021). These codecs compress audio into discrete codes optimized for reconstruction and are widely used as both inputs and outputs in recent speech language models (Défossez et al., 2024; Hu et al., 2025).
The codec models represent each audio frame using multiple discrete codes, drawn from multiple codebooks. The choice of quantization determines how these codes are modelled. Dependent codebooks, typically produced by residual vector quantization (Zeghidour et al., 2021, RVQ), require hierarchical prediction: the language model predicts the first code, while additional modules generate the remaining layers (Défossez et al., 2024; Wang et al., 2023). In this work, we choose independent codebooks from finite scalar quantization (Mentzer et al., 2024, FSQ), which simplify modelling by allowing all codes for a frame to be predicted jointly (Hu et al., 2025). Additional details on encoding and reconstruction are provided in App.˜A.
@@ 3 Related Work
End-to-end architectures typically support these behaviours natively and have therefore become increasingly popular. Among them, Moshi (Défossez et al., 2024) and its variants (Ohashi et al., 2025; Shi et al., 2025) represent a common design choice: employing separate system, user, and text streams, and codec-based models for speech input and output (Défossez et al., 2024, 2022; Zeghidour et al., 2021). Other approaches use continuous features at the input (Hu et al., 2025; Fu et al., 2025), or at both input and output (Yu et al., 2024). Some architectures avoid multiple input–output streams by interleaving modalities (Veluri et al., 2024; Lee et al., 2025; Yu et al., 2024), or by embedding them jointly at the token level (Wang et al., 2025). For more details, we refer the reader to Arora et al. (2025a).
@@ 4 Model
F-Actor, our full-duplex model, is based around an LLM backbone, augmented with a speaker embedding model and an audio encoder. Unlike previous work (Défossez et al., 2024; Hu et al., 2025), F-Actor is designed with instruction following in mind, that is, incorporating both a speaker embedding and an instruction-following prompt. In the following, we describe each component of the system in detail. An overview is provided in Fig.˜1.
@@
NanoCodec has two advantages: (i) it uses FSQ quantization (Mentzer et al., 2024) so the four codebooks are independent, enabling parallel prediction without requiring a depth-transformer architecture (Défossez et al., 2024), and (ii) it has shown strong performance on speech synthesis (Hu et al., 2025).
@@
Text Stream. Following Hu et al. (2025); Défossez et al. (2024), we not only generate audio, but also experiment with generating the corresponding text, using an additional system-side text stream. To do so, we need to align the text tokens and audio DAUs (alignment strategies are described in Section˜6). After embedding the text stream, we add it to the system’s speech stream. The LLM uses this summed embedding as input. We then use the LLM’s original language-modeling head to predict the next text token.
@@ 6.1 Data
Text-Speech Alignment. As described in Section˜4, we experiment with generating not only speech, but also the corresponding text. We explore both word- and utterance-level alignment of speech and text, as prior work has shown mixed results (Défossez et al., 2024; Hu et al., 2025). To this end, we align text and speech tokens using forced alignment with Kaldi (Povey et al., 2011) and a public model888[Uncaptioned image] kaldi-model-m13 trained on Librispeech Panayotov et al. (2015). We exclude conversations with text–speech mismatches due to occasional TTS deviations from the transcript, resulting in 74.4% of the original training set.
@@ 6.2 Training Details
To maximize efficiency, our approach utilizes a single training stage that avoids additional speech or text pretraining as required by prior models such as Défossez et al. (2024). Only the LLM backbone and the linear DAU heads are trained, while the speech encoder remains frozen. Training is performed on four NVIDIA A100-SXM4-40GB GPUs with early stopping based on validation performance, resulting in around 48 hours of training per model. We use a maximum sequence length of 2,048 tokens, a per-GPU batch size of 1, and 8 gradient accumulation steps. Training and inference parameters are provided in Tab.˜6 in App.˜D.
@@ 6.3 Experiments
Text-Stream. We examine the effect of adding a text stream on the system side (see Section˜4), enabling the model to predict not only audio but also the corresponding text (Défossez et al., 2024; Hu et al., 2025). We further experiment with introducing special tokens that precede BCs and interruptions (BC/I tok.) to help the model generate the numbers specified in the prompt.
@@
Audio Delay. We experiment with delaying the audio stream by up to two tokens relative to the text stream, generating text tokens first while padding the audio stream before audio token generation begins (Défossez et al., 2024; Hu et al., 2025).
RVQ vs. FSQ. All primary experiments use an FSQ encoder to benefit from independent codebooks. To test the generality of our multi-head DAU prediction architecture, we also evaluate an RVQ-based setup using Mimi101010[Uncaptioned image] kyutai/mimi (Défossez et al., 2024) with 8 codebooks. This allows us to verify whether predicting DAUs for each codebook in parallel remains effective with RVQ quantization.
@@ 7.3 Comparing Turn-Taking to SOTA Models
In Tab.˜4, we compare turn-taking behaviour to Moshi (Défossez et al., 2024) and dGSLM (Nguyen et al., 2023), with all metrics reported as cumulative durations per category. For all four metrics, Inter-Pausal Units (IPUs), intra-turn Pause, between-turn Gap and Overlap, our model exhibits behavior similar to the training data and the baseline models. Like Moshi and non-cascaded dGLSM, our model’s pauses are generally longer than the gaps, reflecting real-world conversational statistics (Heldner and Edlund, 2010).
@@
7.4 The Role of Sampling Parameters
Similar to Défossez et al. (2024), we observe that sampling temperature has a significant impact on model behaviour. Results for different temperatures (0.60.61.01.0) using our best-performing model (word-level alignment; audio delay of 2) on the dev set are reported in Tab.˜7 in App.˜E. As temperature increases, the dialogue becomes more dynamic, with more BCs and interruptions, more balanced speaking time between the speakers, and improved coherence and adherence to the narrative. However, higher temperatures also lead to slightly lower UTMOS scores and higher WER, reflecting a mild impact on speech quality. In our setup, a temperature of $0.9$ provides the best trade-off, and all results reported in this paper use this setting.
@@ A.1 Discrete Representations of Speech
The first category of discrete representations fo speech consists of units often termed “acoustic tokens” or Discrete Acoustic Units (DAUs), derived from reconstruction-based Neural Audio Codecs (NACs) such as Mimi, EnCodec, or SoundStream (Défossez et al., 2024, 2022; Zeghidour et al., 2021). These units are optimized to compress and perfectly reconstruct the input signal. Consequently, they encode all acoustic information, including speaker identity, prosody, and recording conditions, into a single stream that can be decoded directly back into a waveform. While this enables end-to-end modeling such as in Moshi and VALL-E (Défossez et al., 2024; Wang et al., 2023), the high information density of DAUs can make modeling stability challenging.
@@ A.2 Quantization Methods
The choice of quantization method fundamentally dictates the architecture of the Speech Language Model, specifically regarding how it models the multiple codes representing a single time step. The two primary paradigms involve dependent codebooks, typically resulting from Residual Vector Quantization (Zeghidour et al., 2021, RVQ), and independent codebooks, such as those derived from Finite Scalar Quantization (Mentzer et al., 2024, FSQ). In RVQ-based models like EnCodec or Mimi (Défossez et al., 2022, 2024), the codebooks operate sequentially where each quantizes the residual error of the previous one. This creates a strong dependency among the codes within a single frame. Therefore, generation often requires a hierarchical approach. The primary language model autoregressively predicts the code for the first RVQ layer, while a secondary mechanism, such as the Depth Transformer in Moshi or the nonautoregressive model in VALL-E, predicts the codes for subsequent RVQ layers for that frame (Défossez et al., 2024; Wang et al., 2023). Conversely, methods like NanoCodec use FSQ to project the latent space into independent subspaces, resulting in codebooks that are statistically independent (Casanova et al., 2025). This independence simplifies the modeling task by allowing the Language Model to predict all codes for a frame simultaneously. Such models, as seen in Hu et al. (2025), typically utilize a multiple head output layer, termed a Hydra head, which predicts the indices for all independent codebooks in parallel without the need for an additional sequential depthwise modeling step.
@@ Appendix D Training Details
For inference, we show that sampling parameters, in particular the temperature, have a great influence on model performance in Section˜7.4. We find that for our setup, a temperature of $0.9$ works best and hence report all results (except for those in Tab.˜7) with this temperature.
@@
Table 8: Instruction-following results for our best model (text-stream, audio delay of 2, word-alignment, trained on system + user) trained with FSQ or RVQ speech codecs.